# FaceLift: Single Image to 3D Head with View Generation and GS-LRM

### v1

@@ 4 Experiments
To evaluate single-image head reconstruction methods, we leverage publicly available multiview head datasets to assess novel view synthesis performance using known camera parameters. We present both qualitative comparisons and quantitative error metrics across these datasets.
4.1 Experimental Setting
Evaluation Datasets. We evaluate our method using two multiview datasets: The synthetic Cafca dataset [6]: While this dataset shares our focus on faces, it differs significantly from our training dataset, which was independently developed. For evaluation, we select 40 subjects with 30 test camera poses each. Since the camera positions are randomly distributed, we manually select the most frontal view as input. The Ava-256 dataset [34]: This studio-captured dataset contains real human subjects. We sample 10 subjects and 10 test camera poses for our evaluation (more details are in the supplementary material). To demonstrate our system’s generalization capabilities, we also evaluate on a collected set of in-the-wild face images for qualitative assessment.
Baselines. We compare our results with three recent works: PanoHead [1], Era3D [26], and LGM [51]. PanoHead [1] is the state-of-the-art method for 3D full head reconstruction using 3D GANs. We perform GAN-inversion to obtain the reconstruction from a single input image. Note that it can only render images within a specific camera pose range and lacks a true 3D representation for novel view synthesis. Era3D [26] is a recent method for general object reconstruction. It employs a multi-view diffusion model to generate six views and uses the Instant-NSR-PL [17] implementation of NeuS [56] for mesh extraction. LGM [51], similar to GS-LRM [64], generates Gaussian splats as a 3D representation after a view generation stage.
Evaluation Metric. We evaluate reconstruction quality using four standard metrics: PSNR, SSIM, LPIPS [65], and DreamSim [12]. To evaluate identity preservation, we perform face verification using ArcFace [10] through the DeepFace [46] implementation.
Implementation Details. Both Cafca [6] and Ava-256 [34] datasets provide multi-view RGB images and corresponding camera poses. However, their camera systems differ from the ones utilized in FaceLift and baselines. We recalculate the test camera extrinsic in each method’s camera system. For a more accurate comparison, we use the Mediapipe facial landmark detection algorithm [33] to identify facial landmarks in both target images and rendered outputs, aligning them based on landmark distributions. Details of this alignment process are provided in the supplementary material. Note that aligning images based on facial landmark detection only works for images that capture the front facial features.
@@ 4.1 Experimental Setting
[…]
Figure 5: Visual results on Cafca dataset [6]. FaceLift provides rendering results that closer align with the ground truth. Notice that PanoHead [1] can not handle the challenging hairstyle in row 1, Era3D [26] creates artifacts on the back of the head, and LGM [51] delivers inaccurate nose and jaw shapes.
4.2 Results on Cafca Dataset
We report numerical comparison results before and after facial landmark alignment in Tab. 1. Before facial landmark alignment, all 30 camera poses in the Cafca dataset are utilized as test views; some views do not cover the face region. Hence, we only report the face identity preservation metric ArcFace [10] in the aligned comparison. For the aligned comparison, each subject has 7 to 19 test views. The test views utilized in the experiment can be found in the supplementary material. In both comparisons, FaceLift performs favorably against baselines, especially for DreamSim [12], which has a better alignment with human similarity judgments. It also achieves a better performance on identity preservation, demonstrated by a lower ArcFace [10] face embedding distance.
We show visual results on the Cafca dataset in Fig. 5. FaceLift renders novel views that better align with the ground truth, while other methods often fail to reconstruct the 3D head in correct colors or geometry structures.
4.3 Results on Ava-256 Dataset
@@
Table 2: Numerical results on Ava-256 dataset [44]. FaceLift performs favorably than baseline methods in both reconstruction metrics and identity facial identity metric, showing a better generalization ability towards real-captured human images.
@@
Figure 6: Visual results on Ava-256 dataset [44]. Compared with baseline methods, FaceLift provides multi-view renderings that are more realistic and similar to ground truth. Era3D fails to deliver delicate facial structure, while LGM generates heads in inaccurate shapes and colors.
@@
We further evaluate FaceLift against other baselines on a studio-captured real human dataset, Ava-256 [44]. PanoHead [1] fails to produce reasonable results with the test camera poses in this dataset, so we exclude this baseline. Tab. 2 shows that FaceLift outperforms all other baselines across all evaluation metrics, demonstrating superior reconstruction quality and identity preservation. It also highlights FaceLift’s strong ability to generalize to real human faces. As shown in Fig. 6, FaceLift achieves more realistic head synthesis, while Era3D [26] struggles with accurate skin and hair textures, as well as facial details. LGM [51] produces inaccuracies in the nose shape.
@@
Figure 7: Visual comparison on in-the-wild data. FaceLift demonstrates great generalization ability and robustness towards in-the-wild images, provides realistic unseen view rendering results. Era3D [26] and LGM [51] generate 3D head representation in inaccurate shape. PanoHead [1] often creates severe artifacts on the back of the head and can not handle challenging hairstyles well.
4.4 Results on In-the-wild Images
To further demonstrate FaceLift’s generalization ability and highlight its potential for real-world applications, we collect in-the-wild human face images and present qualitative results in comparison with other baselines in Fig. 7. Baseline methods often produce undesirable artifacts. For example, PanoHead [1] frequently fails to render the back of the head and occasionally generates extra eyes on the rear. It also struggles with accurately synthesizing hair, shadows, wrinkles, and facial paint. Meanwhile, PanoHead’s outputs lack multi-view consistency, as seen in the last row, where the girl continues to face the camera in novel view 1 despite a change in camera pose. Era3D [26] often struggles to produce an accurate head shape, particularly from the left and right views, and provides fewer geometric details compared to FaceLift. LGM [51] generates Gaussians with inaccurate color and opacity, and lacks a proper understanding of facial geometry, resulting in distorted facial features.
Moreover, we present more FaceLift’s novel view rendering results in Fig. 8 to demonstrate FaceLift’s ability to produce high-fidelity, realistic 3D head reconstructions with intricate details across a variety of challenging scenarios. FaceLift effectively handles faces under various lighting conditions. It can especially render realistic novel view images given a photo captured with an iPhone under dark lighting conditions (row 1 column 1), emphasizing its robustness and potential for real-world application. It reconstructs facial details with high fidelity, especially the wrinkles and folds on the face caused by extreme expression (row 2, row 3). FaceLift also excels at reconstructing challenging textures, such as mustaches and hair. Furthermore, it faithfully reconstructs facial paint, despite such data not being included in our synthetic face dataset, showcasing its strong generalization ability.
@@ 4.4 Results on In-the-wild Images
Figure 8: Results of FaceLift on in-the-wild images. FaceLift accurately reconstructs 3D head models under challenging lighting conditions, achieving high fidelity (row 1). It captures fine facial details such as wrinkles (row 2), mustaches, and individual hairs (row 3). Additionally, it remains robust to complex facial expressions (row 3), various skin tones (row 4), and hair colors (row 5). Furthermore, it can realistically reconstruct facial paint (row 5).
4.5 Application: Video as Input for 4D NVS
As shown in Fig. 9, benefits from the high-fidelity reconstruction ability, FaceLift can generate consistent 3D Gaussian sequence given a video as input to achieve 4D rendering. We process input video frames sequentially, generate one 3D Gaussian representation for each frame, and form a Gaussian sequence. Then, given any timestamp, we can select the corresponding 3D Gaussians from the Gaussian sequence and render from any given pose. Without any bells and whistles to constrain the consistency between different timestamps, FaceLift can render video sequences under novel poses with minimal artifacts.
Combined with 2D face animation methods like LivePortrait [16], FaceLift achieves 3D face animation. Due to its robustness and strong generalizability, FaceLift can lift faces with extreme facial expressions to 3D without sacrificing fidelity. The results of this application are shown on our website.
Figure 9: Video as input for 4D rendering. Given a video as input, FaceLift processes each frame sequentially and generates 3D Gaussian sequence. By rendering at a given camera pose and timestamp, FaceLift achieves 4D novel view synthesis.
4.6 Ablation Study
Number Views for GS-LRM Reconstruction. Vanilla GS-LRM [64] uses four views for inference. Instead, we add two more views, front-left and front-right, to further improve the reconstruction of the detailed facial structure. We compare baselines using different numbers of images as GS-LRM input in Fig. 11. With only four views, GS-LRM fails to reconstruct a complete forehead, while with six input views, GS-LRM can reconstruct eyes and eyebrows more smoothly. It also more realistically reconstructs challenging textures like wrinkles on the face and folds in the ears.
Figure 11: Ablation study on number of input views to GS-LRM. Using 6 views provides a more complete reconstruction of the forehead and delivers more accurate details on eyes and the geometry structure of wrinkles and ears.
Effectiveness of GS-LRM Fine-tuning. GS-LRM [64] used in FaceLift is fine-tuned from the GS-LRM checkpoint pre-trained on Objaverse [8, 9] with synthetic human head data. To highlight the importance of training on synthetic human head data, we compare FaceLift’s rendering results with GS-LRM trained on Objaverse and our fine-tuned GS-LRM. Without fine-tuning, GS-LRM struggles to accurately estimate facial depth, resulting in artifacts around the eye region and a flat top of the head due to the absence of Gaussians splatted from above. After fine-tuning, the model learns prior knowledge of head structure, producing smoother and more realistic renderings.
Figure 12: Ablation study on GS-LRM fine-tuning. After fine-tuning, our GS-LRM reconstructor gains a more refined understanding of facial geometry, resulting in improved accuracy for facial features such as eyes, noses, and hair.
5 Conclusions and Future Work
@@
We propose FaceLift, a feed-forward approach that lifts a single facial image to a detailed 3D reconstruction with preserved identity features. FaceLift uses multi-view diffusion to generate unobservable views and employs GS-LRM to reconstruct 3D Gaussian splats, enabling high-quality novel view synthesis. To overcome the difficulty of capturing real-world multi-view human head images, we render high-quality synthetic data for training and show that, despite being trained solely on synthetic data, FaceLift can reconstruct 3D heads from real-world captured images with high fidelity. Compared with baselines [1, 26, 51, 63], FaceLift generates 3D head representations with finer geometry and texture details and exhibits better identity preservation ability.
Furthermore, we demonstrate that FaceLift supports 4D novel view synthesis given a video sequence as input, demonstrating our method’s high-fidelity results and robustness. Nevertheless, we still observe some inconsistencies across time stamps, especially with hair. To enhance temporal consistency, we plan to incorporate a 4D Gaussian representation in future work, allowing us to better constrain the position, shape, size, rotation, etc. of the 3D Gaussians across frames.
@@ Appendix A Overview
This supplementary material presents additional results to complement the main manuscript. We provide further experiments in Sec. B, including a comparison with DimensionX [49] and additional visual results of FaceLift on in-the-wild images. We illustrates experimental details in Sec. C. Finally, we discuss the limitations of FaceLift in Sec. D.
Appendix B Additional Experiments
B.1 Comparison with DimensionX
@@
Figure 13: Visual comparison with DimensionX. DimensionX frequently produces inaccuracies in the back of the head and the shoulder shapes. Other common issues include misaligned ears and eyes gazing in incorrect directions. Additionally, controlling camera poses is challenging. In contrast, FaceLift delivers results that are significantly more consistent across multiple views while enabling the generation of more visually appealing hair.
@@
We provide comparison results on single image to 3D tasks with a state-of-the-art video diffusion model, DimensionX [49]. DimensionX is a framework designed to generate photorealistic 3D and 4D scenes from a single image with video diffusion. The results are shown in Fig. 13. As a video diffusion model, DimensionX struggles to produce multi-view consistent results and lacks a clear spatial understanding of head shapes. As a result, it often generates eyes gazing in the wrong direction and ears positioned incorrectly, along with inaccurate shoulder shapes. In contrast, FaceLift generates highly realistic 3D human heads while also producing more visually striking hair.
B.2 Additional Results on In-the-wild Images

### v2

@@ 4 Experiments
4.1 Experimental Setup
Evaluation Datasets. To quantitatively evaluate FaceLift, we establish two benchmarks for single-image to 3D full head reconstruction tasks using publicly available datasets: (1) Cafca dataset [8]: We select 40 subjects with 7 to 19 test camera poses each. Since the camera positions are randomly distributed, we manually select the most frontal view as input. Note that this synthetic dataset was independently developed and differs significantly from our training dataset. (2) Ava-256 dataset [39]: This studio-captured dataset contains real human subjects. We sample 10 subjects and 10 test camera poses for our evaluation. More details in supplemental. To demonstrate our system’s generalization capabilities, we also evaluate on a set of in-the-wild face images for qualitative assessment.
Baselines. We compare our results against three state-of-the-art methods for single-face 3D reconstruction: GGHead [29], PanoHead [1], and Dual Encoder [4]. We perform GAN-inversion to reconstruct 3D head from a given input image using these models. To emphasize the importance of utilizing our synthetic human head data for training, we also compare our method with two methods that focus on general object reconstruction: Era3D [31] and LGM [56]. More comparison results with mesh-based methods are provided in the supplementary material.
We further developed a baseline, Our MV + LGM, which leverages the multi-view outputs generated by our diffusion model and employs LGM for reconstruction. This demonstrates that our method can be seamlessly integrated with other reconstruction frameworks to enhance performance on face reconstruction tasks. We tried to fine-tune the LGM reconstructor with our synthetic data, but it provides inferior results with incorrect geometry and artifacts compared with the original weights, which we suspect is due to training data mismatch (see details in the supplementary material).
Evaluation Metrics. We evaluate reconstruction quality using four standard metrics: PSNR, SSIM, LPIPS [75], and DreamSim [16]. To evaluate identity preservation, we perform face verification using ArcFace [14] through the DeepFace [52] implementation.
Implementation Details. Both Cafca [8] and Ava-256 [39] datasets provide multi-view RGB images and corresponding camera poses. However, their camera system differs from the ones utilized in FaceLift and baselines. We recalculate the test camera extrinsic in each method’s camera system. For a more accurate comparison, we use the Mediapipe facial landmark detection algorithm [38] to identify facial landmarks in both the target images and the rendered outputs, aligning them based on landmark distributions. Details are provided in the supplementary material.
Our system takes approximately 8 seconds to infer a 3D Gaussian head from a single image: about 1.5 seconds for preprocessing (background removal, rescaling, etc.), 5.5 seconds for multi-view image generation, and under 1 second for 3D Gaussians reconstruction.
@@ 4.1 Experimental Setup
Table 1: Quantitative results on Cafca. FaceLift achieves favorable performance on all evaluation metrics.
Figure 5: Visual results on Cafca compared with face reconstruction methods. FaceLift renders novel views that closely match the ground truth, while other baselines often fail to reconstruct the 3D head in correct colors or geometry structures.
Figure 6: Visual results on Cafca compared with general objects reconstruction methods. Comparison with general object reconstruction methods shows the importance of specialized data.
4.2 Experiments on the Cafca Dataset
We report numerical comparison results on Cafca in Tab. 1. FaceLift performs favorably against baselines, especially on DreamSim [16] metric, indicating high-quality perceptual similarity. It also achieves better identity preservation performance, as demonstrated by a lower ArcFace [14] distance. We show visual results in Fig. 5 and Fig. 6. FaceLift yields rendering results that closely match the ground truth. Compared with other baselines, GGHead [29] does not support full-head rendering, resulting in unrealistic outputs when the view angle significantly deviates from the input. PanoHead [1] struggles with challenging hairstyles, while Dual Encoder [4] produces blurred facial textures. Additionally, Era3D [31] introduces artifacts on the back of the head, and LGM [56] yields inaccurate nose and jaw shapes, underscoring the importance of our synthetic human head data. When integrated with our multi-view diffusion approach, LGM achieves enhanced performance, demonstrating that our method can be seamlessly combined with existing baselines to boost their results.
@@ 4.2 Experiments on the Cafca Dataset
Table 2: Quantitative results on Ava-256. FaceLift performs favorably than baseline methods in both reconstruction metrics and identity facial identity metric, showing a better generalization ability towards real-captured human images.
@@
Figure 7: Visual results on Ava-256. Compared with baselines, FaceLift provides multi-view renderings that are more realistic and similar to ground truth. Era3D fails to deliver delicate facial structures, while LGM generates inaccurate head shapes and colors.
4.3 Experiments on the Ava-256 Dataset
@@
We further evaluate FaceLift against other baselines on a studio-captured real human dataset, Ava-256 [50]. GAN-inversion based methods [1, 29, 4] fail to produce reasonable results with the test camera poses in this dataset, so we exclude these baselines. Tab. 2 shows that FaceLift outperforms all other baselines across all evaluation metrics, demonstrating superior reconstruction quality and identity preservation. It also highlights FaceLift’s strong ability to generalize to real human faces. As shown in Fig. 7, FaceLift achieves more realistic head synthesis, while Era3D [31] struggles with accurate skin and hair textures, as well as facial details. LGM [56] produces inaccuracies in the nose shape. When combined with our multi-view diffusion model, LGM yields more accurate geometric structures, yet its texture quality remains inferior to that of FaceLift.
@@
Figure 8: Visual comparison on in-the-wild data. FaceLift demonstrates great generalization ability and robustness towards in-the-wild images, provides realistic unseen view rendering results. Era3D [31] and LGM [56] generate 3D head representation in inaccurate shape. PanoHead [1] often creates severe artifacts on the back of the head and can not handle challenging hairstyles well. Dual Encoder [4] shows improved performance on reconstructing the back of the head but exhibits more pronounced identity loss.
@@
Figure 9: Results of FaceLift on in-the-wild images. FaceLift accurately reconstructs 3D head models under challenging lighting conditions, achieving high fidelity (row 1). It captures fine facial details such as wrinkles (row 2), mustaches (row 3), and individual hairs (row 2 and row 4). Additionally, it remains robust to complex facial expressions (row 3) and various skin tones (row 4). Furthermore, it can realistically reconstruct facial paint (row 4). More results are provided in the supplementary materials.
4.4 Experiments with In-the-wild Images
We collect in-the-wild human face images and present qualitative results in comparison with other baselines in Fig. 8. Baseline methods often produce undesirable artifacts. For instance, PanoHead [1] frequently fails to render the back of the head and sometimes generates extra eyes at the rear. It also struggles to synthesize hair, shadows, wrinkles, and facial paint accurately, and its outputs lack multi-view consistency (e.g., the girl continues to face the camera in novel view 1 despite a changed pose). Dual Encoder [4] improves back-of-head rendering but suffers from severe identity loss (row 2) and fails to accurately reconstruct face paint (row 4). Era3D [31] often produces an inaccurate head shape, particularly from side views, and offers fewer geometric details compared to FaceLift. LGM [56] generates Gaussians with inaccurate color and opacity and lacks proper facial geometry, resulting in distorted features. Baseline Our MV + LGM shows that our multi-view diffusion model enhances LGM by providing improved facial geometry and texture details. However, the LGM reconstructor still produces Gaussians with inaccurate shapes and opacities, further underscoring the advantages of our transformer-based reconstructor.
We present more FaceLift’s novel view rendering results in Fig. 9 to demonstrate FaceLift’s ability to produce high-fidelity, realistic 3D head reconstructions with intricate details across a variety of challenging scenarios. FaceLift effectively handles faces under various lighting conditions. It can especially render realistic novel view images given a photo captured with an iPhone under dark lighting conditions (row 1 column 1), emphasizing its robustness and potential for real-world application. It reconstructs facial details with high fidelity, especially the wrinkles and folds on the face caused by extreme expression. FaceLift also excels at reconstructing challenging textures, such as mustaches and hair. Furthermore, it faithfully reconstructs facial paint, despite such data not being included in our synthetic face dataset, showcasing its strong generalization ability.
5 Ablation Study
Figure 10: Importance of input view reconstruction. The diffusion model that is not trained to perform the input view reconstruction, i.e., w/o. Input View Reconstruction, overfits to synthetic training distribution, suffers from severe identity loss during inference. Trained with input view reconstruction, our method preserves the input identity and expression faithfully.
5.1 Input View Reconstruction
We conduct an ablation study to demonstrate the importance of reconstructing the input view during training. For comparison, we train a multi-view diffusion model that generates six novel views. In this baseline, the first generated view’s elevation is adjusted from $0^{\circ}$ to $20^{\circ}$, while the remaining views adopt the same camera poses as in our default setting. We refer to this variant as w/o. Input View Reconstruction. Fig. 10 presents the view generation results of the two diffusion models when applied to real-world images. Without the input view reconstruction task, the model trained on the synthetic dataset generates views within a limited distribution, leading to noticeable identity loss. Moreover, it loses its ability to preserve facial expressions and face paint. In contrast, incorporating the input view reconstruction task during training enables our diffusion model to faithfully regenerate the input view, significantly improving its generalization ability. Quantitative comparison is provided in the supplementary material.
5.2 Number of Views
We evaluate three configurations: four views (front, left, right, and back), six views (adding front-left and front-right), and eight views (further including front-top and front-bottom). Fig. 11 compares the baselines using different numbers of input views. With only four views, the reconstructor fails to capture a complete forehead; however, with six views, it reconstructs the eyes and eyebrows more smoothly and renders challenging textures—such as facial wrinkles and ear folds—more realistically. Eight views do not offer significant visual improvements, and incur a higher computational cost in both stages. Thus, we find that six views achieve a good balance between reconstruction quality and computational efficiency.
Figure 11: Number of input views of Gaussian reconstructor. Using six views strike a good balance between reconstruction quality and computational efficiency.
Figure 12: Two-stage training of reconstructor. Without pre-training on general objects, the reconstructor fails to produce clear textures in the reconstruction results. Meanwhile, without fine-tuning on synthetic human head data, the model lacks a refined understanding of facial structures, including the eyes and nose.
5.3 Two-stage Reconstructor Training
As illustrated in Sec. 3.3, our Gaussian reconstructor follows a two-stage training pipeline. Fig. 12 shows that pre-training on general objects helps the model learn a diverse prior of geometric structures, resulting in clearer textures on delicate facial regions. Meanwhile, fine-tuning on synthetic human head data enables the reconstructor to gain a more refined understanding of facial structure, thereby enhancing the accuracy of features such as the eyes, nose, and hair.
6 Conclusions
@@
We propose FaceLift, a feed-forward approach that lifts a single facial image to a detailed 3D reconstruction with preserved identity features. Our method uses multi-view diffusion to generate unobservable views and employs a transformer-based reconstructor to reconstruct 3D Gaussian splats, enabling high-quality novel view synthesis. To overcome the difficulty of capturing real-world multi-view human head images, we render high-quality synthetic data for training and show that, despite being trained solely on synthetic data, FaceLift can reconstruct 3D heads from real-world captured images with high fidelity. Compared with baselines [1, 4, 29, 31, 56], FaceLift generates 3D head representation with finer geometry and texture details and exhibits better identity preservation ability.
@@ 7 Acknowledgement
This was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).
@@ 1 Overview
This supplementary material presents additional results to complement the main manuscript. We first provide a supplementary video showcasing additional visual results. We then provide further experiments in Sec. 3, including a comparison with DimensionX [55], additional visual results of FaceLift on in-the-wild images, additional ablation study results and an autoregressive generation pipeline to apply FaceLift on videos to achieve 4D rendering. We deliver more details on our method in Sec. 4 and illustrate experimental details in Sec. 5. Finally, we discuss the limitations of FaceLift in Sec. 6.
2 Supplementary Video
Please refer to our supplementary video for a more comprehensive visualization of the results. The video includes additional examples of single-image-to-3D head reconstruction, demonstrations in the interactive viewer, and results showcasing video-based input for 4D novel view synthesis.
3 Additional Experiments
3.1 Comparison with DimensionX
@@
Figure 13: Visual comparison with DimensionX [55]. DimensionX frequently produces inaccuracies in the back of the head and the shoulder shapes. Other common issues include misaligned ears and eyes gazing in incorrect directions. Additionally, controlling camera poses is challenging. In contrast, FaceLift delivers results that are significantly more consistent across multiple views while enabling the generation of more visually appealing hair.
@@
We provide additional comparison results on single image to 3D tasks with a state-of-the-art video diffusion model, DimensionX [55]. DimensionX is a framework designed to generate photorealistic 3D and 4D scenes from a single image with video diffusion. The results are shown in Fig. 13. As a video diffusion model, DimensionX struggles to produce multi-view consistent results and lacks a clear spatial understanding of head shapes. As a result, it often generates eyes gazing in the wrong direction and ears positioned incorrectly, along with inaccurate shoulder shapes. In contrast, FaceLift generates highly realistic 3D human heads while also producing more visually striking hair.
3.2 Comparison with Mesh-based Methods
We provide comparison results with mesh-based reconstruction methods InstantMesh [70], Unique3D [67], and TRELLIS [69] on the Cafca dataset [8]. Quantitative results are shown in Tab. 3, and quantitative comparisons are shown in Fig. 14. Results show that mesh-based reconstruction methods fail to provide realistic hair texture and detailed skin wrinkles. Meanwhile, thanks to the input view reconstruction strategy, FaceLift achieves superior identity preservation.
Figure 14: Visual results on Cafca compared with mesh-based reconstruction methods. Compared to mesh-based reconstruction methods, our use of pixel-aligned 3D Gaussians offers clear advantages: the semi-transparent kernels naturally capture complex visual phenomena such as hair strands and fine wrinkles.
TRELLIS [69] 12.74 0.7412 0.3746 0.2170 0.4001
Unique3D [67] 14.27 0.7643 0.3188 0.1277 0.2088
InstantMesh [70] 16.44 0.7815 0.2792 0.1504 0.2741
FaceLift 16.61 0.7968 0.2694 0.1096 0.1573
Table 3: Quantitative results on Cafca compared with mesh-based reconstruction methods. FaceLift achieves better quantitative results with more suitable representations and specialized training data.
3.3 Additional Results on In-the-wild Images